OpenStudy (quickstudent):

Is this correct? question: How do you find the area of a circle? my answer: The area of a circle can be found using this formula: A = πr^2, where r is the radius

OpenStudy (quickstudent):

@mathstudent55

OpenStudy (quickstudent):

@pooja195

OpenStudy (chrishogan):

Hello, I can hep you. :)

OpenStudy (chrishogan):

Help**

OpenStudy (chrishogan):

Area equals Pi times r times r (or A = πr^2) Area = 3.14 · (3 in) · (3 in) Area = 3.14 · (9 in) Area = 28.26 in (This is assuming r=3 for the example) So, yes! You're correct. :)

OpenStudy (quickstudent):

Thanks

OpenStudy (chrishogan):

Let me know if you're in need of anymore help. :)